kevin invested part of his $10,000 bonus in a certificate of deposit that paid 6% annual simple interest, and the remainder in a mutual fund that paid 11% simple interest. If his total interest for the year was $900.00, how much did kevin invest in the mutual fund?
Answer by richwmiller(17219)

We know the total amount of money invested. $10000
x+y=10000,
We know that the total interest for the year for the two accounts is $900
0.06*x+0.11*y=900
x=10000-y
We substitute for x
0.06*(10000-y)+0.11*y=900
We multiply out
600-0.06y+0.11*y=900
We combine like terms.
0.05*y=300
Isolate y
y=6000 at 11%
x=10000-y
Calculate x
x=4000 at 6%
We check
0.06*4000+0.11*6000=900
240+660=900
900=900
Since this statement is TRUE and neither amount is negative then it is ok